OpenStudy (riley):

Find the derivative of the function. f(x) = −x + 2/x +1

OpenStudy (riley):

I think I figured it out, but I'm not sure. >_<

OpenStudy (riley):

the -x and +1 both get dropped.

OpenStudy (riley):

so wouldn't that just leave -2/x^2

OpenStudy (anonymous):

=-1 - 2/(x+1)^2

OpenStudy (anonymous):

I'm not of the last term in f(x) = −x + 2/x +1 is in the numerator or not

OpenStudy (riley):

oh wait sorry, I should of wrote it more clearly. the problem is: f(x) = -x + (2/x) +1

OpenStudy (anonymous):

if not then f'(x) = -1-2/x^2

OpenStudy (riley):

so I was right, then it's -2/x^2!!??

OpenStudy (anonymous):

it's ok now.so your answer is correct.

OpenStudy (riley):

yay, can't believe i did that right lol. thank you so very much ^_^